# A Survey on Quantum Machine Learning: Basics, Current Trends, Challenges, Opportunities, and the Road Ahead

### v2

@@ IV-D Quantum Neural Networks
What makes QNNs exciting compared to traditional ANNs is the differentiable nature of the quantum circuits. Quantum computers can compute the changes in the control parameters needed to make the QNN better at a given task. While common applications of QNNs are in the area of image recognition and object detection since they excel in detecting specific features from images, QNNs can be used for a plethora of applications, as discussed in Section VII.
@@
IV-D1 Quantum Convolutional Neural Networks
Convolutional Neural Networks(CNNs) are established architectures for image classification tasks in the computer vision domain [88]. Quantum Convolutional Neural Networks (QCNNs) are HQNN architectures whose structure is inspired by classical CNNs (see Figure 22). The QCNN circuit model proposed by Cong et al. [89] extends fundamental properties of CNNs to the quantum domain by using only $O(log(N))$ variational parameters for input sizes of $N$ qubits. Therefore, it enables efficient training and implementation of QCNNs on realistic, near-term quantum devices.
@@ IV-D1 Quantum Convolutional Neural Networks
While the classification task is the primary purpose for using QCNNs, they can also devise a quantum error correction scheme optimized for a given error model [89]. However, QCNNs with large circuit depths are challenging to implement due to the high coherence required for qubits.
IV-D2 Quanvolutional Neural Networks
Motivated by the idea of a convolution layer where, instead of processing the entire input data with a global function, a local convolution is applied, Quanvolutional Neural Networks (QuanCNNs) [90] are based on a “quanvolutional” kernel. Compared to QCNNs, where the complete architecture is implemented with PQCs, QuanCNNs focus only on building efficient convolutional layers using quantum circuits (see Figure 23). Considering that the underlying basic principle of convolutional layers is to extract features from data hierarchically, a quanvolutional layer leverages the ability of quantum systems to contextualize greater information capacity. Therefore, QuanCNNs introduce preliminary layers before the classical CNN to manipulate (i.e., pre-process through transformations) the inputs using a single or a set of consecutive quantum layers. The input of the first classical layer is a quantum output measured classically that preserves a large amount of information for improving the contextual feature extraction in the classical layers. In other words, it aims to encapsulate the input information using a representation that results from the quantum space transformations. This approach complements and facilitates the architecture in identifying correlated features in the subsequent (classical) layers.
@@ IV-D2 Quanvolutional Neural Networks
IV-D3 Qauntum Recurrent Neural Networks
Recurrent Neural Networks (RNNs) are the foundational models for using sequential data to solve temporal problems such as natural language processing. Inspired by the structure of a VQE quantum circuit, a Quantum Recurrent Neural Network (QRNN) replaces the RNN layers with PQCs. While VQE circuits are very dense, QRNN cells are highly structured circuits with fewer parameters that are reused. Each parameter has a higher-level logical unit than the VQE components [91]. Applying multiple QRNN cells iteratively on the input sequence generates a QRNN (see Figure 24) that behaves similarly to classical RNNs. However, due to the decoherence issues, the QRNNs’ performance on long sequential data is typically worse than classical RNNs. The coherence constraints can be relaxed if the QRNN architecture is built by stacking the recurrent blocks in a staggered way [92].
@@ IV-D3 Qauntum Recurrent Neural Networks
IV-D4 Quantum Generative Adversarial Networks
A Generative Adversarial Network (GAN) aims to generate data that resembles the original samples used in the training dataset. This is typically implemented using a generator that produces new (fake) data and a discriminator that distinguishes fake from real data. Fully Quantum GANs (QGANs) typically require too large resources that cannot be implemented in near-term quantum devices. To overcome this limitation, it is possible to design hybrid QGANs where either the generator or the discriminator is implemented with classical computing. For example, the QGAN architecture proposed by Huang et al. [93] has a quantum generator and a classical discriminator. The generator is divided into multiple sub-generator blocks, each of them responsible for generating a patch of the full image. As shown in Figure 25, the generated image is formed by concatenating the patches. This approach can be implemented on quantum computers with a limited number of available qubits, where the same device can execute the sub-generator sequentially. This architecture provides guidance for developing advanced QGAN models on near-term quantum devices and opens up an avenue for exploring quantum advantages in various GAN-related tasks.
@@ IV-D4 Quantum Generative Adversarial Networks
IV-D5 Quantum Autoencoders
Autoencoders (AEs) are self-supervised ML models that reduce the size of the input data by reconstructing it. It can compress and encode information from the input using representation learning. Quantum counterparts of AEs, called Quantum Autoencoders (QAEs) [94], aim to reduce the dimensionality of quantum states. As shown in Figure 26, a QAE is composed of three layers, which are input, bottleneck, and output. Its primary use is for digital compression, where the information can be encoded into a smaller amount of qubits. QAEs enable the transformation and mapping of large problems into smaller quantum circuit equivalents. Moreover, QAEs are useful for denoising since they extract relevant features from the initial quantum state into encoded data while neglecting the additional noise [50].
@@ IV-D5 Quantum Autoencoders
IV-D6 Binary Neural Networks with Quantum Optimization
Another class of neural networks, whose design is motivated by the need for a compact architecture for increased portability and ubiquitous computing, is the Binary Neural Networks (BNNs). As suggested by their name, the working principle is to replace floating point feature weights and activations with binary values, introducing a binarization process. The redesigned strategy is favorable as it reduces the required storage and computation for resource-constrained devices while leveraging the trade-off with its performance. This process allows compact, lightweight, and power-efficient networks that can simultaneously maintain acceptable accuracy. Methods for improving the performance of BNNs are an active area of research and quantum methods theoretically present themselves as a useful tool for improvement. The work presented in [95] argued theoretically that QCs could train BNNs to their global optimum via quantum search techniques in a shorter time than using brute-force classical optimization. An alternative approach to training BNNs utilizes a quantum superposition of weights [96, 97].
IV-D7 Single-Qubit Neural Networks
Single-qubit QNNs are based on the data re-uploading technique [98]. According to this method, each layer receives a copy of the input on the same single qubit line. The measured state of the solution stores the sequentially-encoded processing result carried on from all the sequential layers. The work in [99] presents a theoretical framework formulated to showcase the expressive capability of the employed data re-uploading technique over the quantum neural network layers, each of them comprising interconnected encoding circuits and trainable circuit blocks. It is further highlighted how single-qubit QNNs can approximate any univariate function by mapping the model to a partial Fourier series.While it seems to be a useful direction to be explored, as demonstrated in [100], there are limitations for evaluating multivariate functions by analyzing the frequency spectrum and the flexibility of Fourier coefficients which must be kept in mind while designing architectures based on the single qubit QNN template.  [99, 101, 102]
IV-E Quantum Kernels
@@
Constructing QKs for SVMs has been under investigation in recent works. The performance of the resulting quantum SVMs for classification problems with a small number of training points exceeds that of optimized classical models with conventional kernels [103][104]. Moreover, QKs allow us to interpret the learning process with the help of quantum information tools [105].
Quantum clustering is a class of data clustering techniques that use the mathematical and conceptual tools of quantum mechanics. It belongs to a family of density-based clustering algorithms, where clusters are defined by areas of higher-density data points. Its design involves a scale-space probability function viewed at the lowest eigenstate of a Schrödinger equation, followed by simple analytic operations to derive a potential function whose minima determine cluster centers (see Figure 27). The design proposed by Horn et al. [106] was applied to a two-dimensional system but is also scalable to higher dimensions. As demonstrated by Aïmeur et al. [107], quantum clustering for different algorithms (divisive clustering, k-medians, and construction of a neighborhood graph) can obtain significant speedup compared to classical clustering.
@@
IV-F Quantum Computing for Bayesian Machine Learning
Bayesian ML is based on utilizing Bayesian methods, like Gaussian processes, to carry out ML tasks. Providing information on the uncertainty of predictions is one of its greatest strengths [108]. Looking into channeling this aspect of advantage to creatively encompass the uncertainty in a quantum methodology is a promising research direction. Interesting results have been obtained connecting deep feed-forward neural networks with Gaussian processes, which eliminates the need for back-propagation while training [108]. As back-propagation itself represents one of the challenges in QML, there is great potential for the success of methods such as Bayesian ML that can perform ML without back-propagation. This can be further optimized by utilizing quantum components in other useful paradigms within the framework. Specialized algorithms can find their way to implementation based on the Bayesian ML technique inspired by quantum optimizations. Such algorithms effectively exploit the connection between ML and Gaussian processes from the quantum lens to curate quantum-enhanced frameworks [108].
IV-G Limitations of Current QML Approaches
@@
  • For wide QNNs, the well-known Barren plateau problem [109] leads to trainability issues.
@@ V-C Data Pre-Processing
Since QML has not yet extensively demonstrated its advantages compared to classical ML, it is more susceptible to adoption by the industry. The current quantum computers have few qubits to test and are noisy, making it difficult to demonstrate the current and potential quantum advantage of QML methods. Utilizing quantum methods for ML in the pre-processing phase, it is possible to achieve better classical encoding and performance of quantum classifiers [110]. Since many proposed QML applications rely on using well-known datasets, especially in the context of quantum data with constraints generated due to the underlying quantum mechanics for representing it as quantum states, it is important to formulate and standardize its collection and encoding processes over time to maintain quality [111].
V-D Data Encoding
Even though the current quantum resources and methodologies provide close estimations and approximations of real-world systems, they do not always constitute adequate datasets for quantum classification models. The most common approach in the NISQ era is transforming the well-known classical datasets into the corresponding quantum representation and vice versa [111]. The fundamental step in the quantum processing pipeline is referred to as state preparation. It allows us to process data on quantum circuits before applying any specific quantum computation. Figure 27 shows an example of how quantum encoding is used to project data from the classical space to the quantum (Hilbert) space.
For tackling the data constraints in QML implementations, data encoding is a fundamental step for representing classical data as quantum states. Encoding layers largely influence the QML model expressivity since the data encoding strategy defines and drives the relevant QML parameters, e.g., the features the QML model can represent [25]. Knowing the various encoding techniques is necessary to choose the most suitable one for solving the particular QML task. Each encoding method has constraints due to the unstable quantum mechanical properties that hinder the complete data encapsulation in quantum representation. However, despite that, they successfully encode the information into quantum states and have large margins of improvement over time [112, 113, 114]. Figure 29 shows a generic overview of the quantum state preparation stage. The process of embedding/encoding generates a quantum representation of the classical data governed by the selected encoding technique within the Hilbert space dimension.
@@ V-D Data Encoding
    The Basis Embedding encodes the binary feature vector into a basis state. It is primarily used when real numbers are mathematically manipulated in quantum algorithms. Such an encoding represents real numbers as binary numbers and transforms them into quantum states. However, since the binary features are not differentiable, the basis encoding does not allow gradient computations [115].
@@
    In the Amplitude Embedding technique, classical data is encoded into the amplitudes of a quantum state. Besides defining the measurement probabilities, each quantum system’s wavefunction can be used to represent its amplitude as a data value. The main advantage of the amplitude embedding is that it can encode $n$ real values (or $n$ fixed-point precision approximation of real numbers) using $\mathcal{O}(\log{n})$ qubits, i.e., it has a logarithmic runtime dependency on the dataset size. Due to its efficiency, it is employed in many QML algorithms [115].
@@
    The Angle Embedding is performed by applying rotations on the x-axis or y-axis using specialized quantum gates, called Pauli rotation gates, along with the values that must be encoded. If we apply angle embedding on a dataset, the number of rotations is the same as the number of features in the dataset [115].
@@
    The Quantum Random Access Memory (QRAM) Encoding [116] is a mechanism that allows accessing classically stored information in superposition by querying an index register. In other words, it enables direct access to the quantum data given the corresponding classical data value. By introducing a conditional rotation and branch selection procedure, we can feed the QRAM with the classical address of the required quantum data and get the qubit state at the output.
@@
    The Qsample Encoding associates a real amplitude vector with a classical discrete probability distribution. It can be considered a hybrid case of Basis and Amplitude Encoding because amplitudes represent the encoded information, but the features are encoded in the qubits [115].
@@
    The Instantaneous Quantum Polynomial-time (IQP) Embedding [117] employs the so-called IQP circuit, which is a quantum circuit of a block of Hadamard gates, followed by a set of gates that are diagonal in the computational basis. Each diagonal gate is composed of a single-qubit RZ rotation gate, encoding the $n$ features, followed by a two-qubit ZZ entangler. The entangler, whose pattern can be customized, encodes the product between features.
    TABLE II: Quantum datasets with key features.
@@
    QDataSet [118] One-qubit and two-qubit systems evolving $14$ TB (compressed), $10,000$ samples each Aug.2021
    QM7, QM8, QM9 [119], [120] [121], [122] Organic molecules Between $7,165$ and $133,885$ molecules, with $7$-$9$ heavy atoms 2012-2015
    PennyLane Datasets [48] Organic molecules and spin $42$ geometries of $34$ molecules, $4$ spin models with $100$ configurations Apr.2022
    TensorFlow Quantum Data [123] Image classification One-qubit per pixel Mar.2020
    Mendley Datasets for Quantum Circuit Mapping [124] Circuit mapping problem as classification task $188,434$ random quantum circuits Sep.2021
    QFlow Lite [125] Charge state recognition $1,001$ $1,599$ measurements, $100\times 100$ $250\times 250$ pixel maps Sep.2018
    NTangled Datasets [126] Quantum states with multipartite entanglement $12,040$ training samples, $4,060$ test samples Nov.2021
@@
    The Quantum Approximate Optimization Algorithm (QAOA) Embedding [80] encodes $n$ features into $m$ qubits, where $m>n$. It uses a layered trainable quantum circuit inspired by the QAOA Ansatz. The feature-encoding circuit associates features with the angles of RX rotations. The main advantage is that it supports gradient computations with respect to both the features and the weights.
@@
    The Hamiltonian Embedding scheme encodes features into the values of the Hamiltonian operator matrix [115]. The Hamiltonian operator of a system is a quantum mechanical operator that corresponds to the total energy of the system. Unlike previous methods that encode the data in the qubits’ states, the Hamiltonian embedding method encodes the data into the operator. Since it is highly correlated with physics, various physical QML algorithms, including VQE and QAOA, are developed using the Hamiltonian encoding method.
@@
    The Displacement Embedding encodes $n$ features into the displacement of amplitudes or phases of $m$ modes, where $n\leq m$. It is used in continuous-variable quantum computing models, where classical information is encoded in the displacement operator parameters [25].
V-E Quantum Datasets Collection

### v3

@@ 4.4 Quantum Neural Networks
Quantum Neural Networks (QNNs) are computational Artificial Neural Network (ANN) models that are based on the principles of quantum mechanics (Panella & Martinelli, 2011). As shown in Figure 20, typically in QNNs, the input data is loaded with classical data inputs. The quantum circuit contains a feature map module with input parameters and an Ansatz module with trainable weights. Measurements are conducted to obtain the outputs.
What makes QNNs exciting compared to traditional ANNs is the differentiable nature of the quantum circuits. Quantum computers can compute the changes in the control parameters needed to make the QNN better at a given task. While common applications of QNNs are in the area of image recognition and object detection since they excel in detecting specific features from images, QNNs can be used for a plethora of applications, as discussed in Section 7.
@@
4.4.1 Quantum Convolutional Neural Networks
Convolutional Neural Networks(CNNs) are established architectures for image classification tasks in the computer vision domain (Capra et al., 2020). Quantum Convolutional Neural Networks (QCNNs) are HQNN architectures whose structure is inspired by classical CNNs (see Figure 23). The QCNN circuit model proposed by Cong et al. (Cong et al., 2019) extends fundamental properties of CNNs to the quantum domain by using only $O(log(N))$ variational parameters for input sizes of $N$ qubits. Therefore, it enables efficient training and implementation of QCNNs on realistic, near-term quantum devices.
@@ 4.4.1 Quantum Convolutional Neural Networks
While the classification task is the primary purpose for using QCNNs, they can also devise a quantum error correction scheme optimized for a given error model (Cong et al., 2019). However, QCNNs with large circuit depths are challenging to implement due to the high coherence required for qubits.
4.4.2 Quanvolutional Neural Networks
Motivated by the idea of a convolution layer where, instead of processing the entire input data with a global function, a local convolution is applied, Quanvolutional Neural Networks (QuanCNNs) (Henderson et al., 2020) are based on a “quanvolutional” kernel. Compared to QCNNs, where the complete architecture is implemented with PQCs, QuanCNNs focus only on building efficient convolutional layers using quantum circuits (see Figure 23). Considering that the underlying basic principle of convolutional layers is to extract features from data hierarchically, a quanvolutional layer leverages the ability of quantum systems to contextualize greater information capacity. Therefore, QuanCNNs introduce preliminary layers before the classical CNN to manipulate (i.e., pre-process through transformations) the inputs using a single or a set of consecutive quantum layers. The input of the first classical layer is a quantum output measured classically that preserves a large amount of information for improving the contextual feature extraction in the classical layers. In other words, it aims to encapsulate the input information using a representation that results from the quantum space transformations. This approach complements and facilitates the architecture in identifying correlated features in the subsequent (classical) layers.
@@ 4.4.2 Quanvolutional Neural Networks
4.4.3 Qauntum Recurrent Neural Networks
Recurrent Neural Networks (RNNs) are the foundational models for using sequential data to solve temporal problems such as natural language processing. Inspired by the structure of a VQE quantum circuit, a Quantum Recurrent Neural Network (QRNN) replaces the RNN layers with PQCs. While VQE circuits are very dense, QRNN cells are highly structured circuits with fewer parameters that are reused. Each parameter has a higher-level logical unit than the VQE components (Bausch, 2020). Applying multiple QRNN cells iteratively on the input sequence generates a QRNN (see Figure 25) that behaves similarly to classical RNNs. However, due to the decoherence issues, the QRNNs’ performance on long sequential data is typically worse than classical RNNs. The coherence constraints can be relaxed if the QRNN architecture is built by stacking the recurrent blocks in a staggered way (Li et al., 2023).
@@ 4.4.3 Qauntum Recurrent Neural Networks
4.4.4 Quantum Generative Adversarial Networks
A Generative Adversarial Network (GAN) aims to generate data that resembles the original samples used in the training dataset. This is typically implemented using a generator that produces new (fake) data and a discriminator that distinguishes fake from real data. Fully Quantum GANs (QGANs) typically require too large resources that cannot be implemented in near-term quantum devices. To overcome this limitation, it is possible to design hybrid QGANs where either the generator or the discriminator is implemented with classical computing. For example, the QGAN architecture proposed by Huang et al. (Huang et al., 2020) has a quantum generator and a classical discriminator. The generator is divided into multiple sub-generator blocks, each of them responsible for generating a patch of the full image. As shown in Figure 25, the generated image is formed by concatenating the patches. This approach can be implemented on quantum computers with a limited number of available qubits, where the same device can execute the sub-generator sequentially. This architecture provides guidance for developing advanced QGAN models on near-term quantum devices and opens up an avenue for exploring quantum advantages in various GAN-related tasks.
4.4.5 Quantum Autoencoders
@@
Autoencoders (AEs) are self-supervised ML models that reduce the size of the input data by reconstructing it. It can compress and encode information from the input using representation learning. Quantum counterparts of AEs, called Quantum Autoencoders (QAEs) (Romero et al., 2017), aim to reduce the dimensionality of quantum states. As shown in Figure 26, a QAE is composed of three layers, which are input, bottleneck, and output. Its primary use is for digital compression, where the information can be encoded into a smaller amount of qubits. QAEs enable the transformation and mapping of large problems into smaller quantum circuit equivalents. Moreover, QAEs are useful for denoising since they extract relevant features from the initial quantum state into encoded data while neglecting the additional noise (Shaib et al., 2021).
4.4.6 Binary Neural Networks with Quantum Optimization
Another class of neural networks, whose design is motivated by the need for a compact architecture for increased portability and ubiquitous computing, is the Binary Neural Networks (BNNs). As suggested by their name, the working principle is to replace floating point feature weights and activations with binary values, introducing a binarization process. The redesigned strategy is favorable as it reduces the required storage and computation for resource-constrained devices while leveraging the trade-off with its performance. This process allows compact, lightweight, and power-efficient networks that can simultaneously maintain acceptable accuracy. Methods for improving the performance of BNNs are an active area of research and quantum methods theoretically present themselves as a useful tool for improvement. The work presented in Liao et al. (2019) argued theoretically that QCs could train BNNs to their global optimum via quantum search techniques in a shorter time than using brute-force classical optimization. An alternative approach to training BNNs utilizes a quantum superposition of weights (Alarcon et al., 2022; Carrasquilla et al., 2023).
4.4.7 Single-Qubit Neural Networks
Single-qubit QNNs are based on the data re-uploading technique (Pérez-Salinas et al., 2020). According to this method, each layer receives a copy of the input on the same single qubit line. The measured state of the solution stores the sequentially-encoded processing result carried on from all the sequential layers. The work in Yu et al. (2022) presents a theoretical framework formulated to showcase the expressive capability of the employed data re-uploading technique over the quantum neural network layers, each of them comprising interconnected encoding circuits and trainable circuit blocks. It is further highlighted how single-qubit QNNs can approximate any univariate function by mapping the model to a partial Fourier series.While it seems to be a useful direction to be explored, as demonstrated in Peña Tapia et al. (2023), there are limitations for evaluating multivariate functions by analyzing the frequency spectrum and the flexibility of Fourier coefficients which must be kept in mind while designing architectures based on the single qubit QNN template (Yu et al., 2022; Goto et al., 2021; Pérez-Salinas et al., 2021).
4.5 Quantum Federated Neural Networks
Figure 27: Functionality of a Quantum Federated Neural Network framework.
The base principle of Federated Learning is to conduct local training and update the global model through global aggregation. In this way, the local users have only access to a subset of the data, to maintain data privacy and high accuracy. Similarly, Quantum Federated Neural Networks (QFNNs) (Innan et al., 2024) use QNNs in an FL environment, demonstrating high accuracy for various applications.
4.6 Quantum Kernels
@@
Constructing QKs for SVMs has been under investigation in recent works. The performance of the resulting quantum SVMs for classification problems with a small number of training points exceeds that of optimized classical models with conventional kernels (Torabian & Krems, 2023; Krunic et al., 2022). Moreover, QKs allow us to interpret the learning process with the help of quantum information tools (Sancho-Lorente et al., 2022).
Quantum clustering is a class of data clustering techniques that use the mathematical and conceptual tools of quantum mechanics. It belongs to a family of density-based clustering algorithms, where clusters are defined by areas of higher-density data points. Its design involves a scale-space probability function viewed at the lowest eigenstate of a Schrödinger equation, followed by simple analytic operations to derive a potential function whose minima determine cluster centers (see Figure 28). The design proposed by Horn & Gottlieb (2002) was applied to a two-dimensional system but is also scalable to higher dimensions.
@@
As demonstrated by Aïmeur et al. (2007), quantum clustering for different algorithms (divisive clustering, k-medians, and construction of a neighborhood graph) can obtain significant speedup compared to classical clustering.
4.7 Quantum Computing for Bayesian Machine Learning
Bayesian ML is based on utilizing Bayesian methods, like Gaussian processes, to carry out ML tasks. Providing information on the uncertainty of predictions is one of its greatest strengths (Zhao et al., 2019). Looking into channeling this aspect of advantage to creatively encompass the uncertainty in a quantum methodology is a promising research direction. Interesting results have been obtained connecting deep feed-forward neural networks with Gaussian processes, which eliminates the need for back-propagation while training (Zhao et al., 2019). As back-propagation itself represents one of the challenges in QML, there is great potential for the success of methods such as Bayesian ML that can perform ML without back-propagation. This can be further optimized by utilizing quantum components in other useful paradigms within the framework. Specialized algorithms can find their way to implementation based on the Bayesian ML technique inspired by quantum optimizations. Such algorithms effectively exploit the connection between ML and Gaussian processes from the quantum lens to curate quantum-enhanced frameworks (Zhao et al., 2019).
4.8 Limitations of Current QML Approaches
@@
  • For wide QNNs, the well-known Barren plateau problem (McClean et al., 2018) leads to trainability issues.
@@ 5.3 Data Pre-Processing
Since QML has not yet extensively demonstrated its advantages compared to classical ML, it is more susceptible to adoption by the industry. The current quantum computers have few qubits to test and are noisy, making it difficult to demonstrate the current and potential quantum advantage of QML methods. Utilizing quantum methods for ML in the pre-processing phase, it is possible to achieve better classical encoding and performance of quantum classifiers (Mancilla & Pere, 2022). Since many proposed QML applications rely on using well-known datasets, especially in the context of quantum data with constraints generated due to the underlying quantum mechanics for representing it as quantum states, it is important to formulate and standardize its collection and encoding processes over time to maintain quality (Sierra-Sosa et al., 2021).
5.4 Data Encoding
Even though the current quantum resources and methodologies provide close estimations and approximations of real-world systems, they do not always constitute adequate datasets for quantum classification models. The most common approach in the NISQ era is transforming the well-known classical datasets into the corresponding quantum representation and vice versa (Sierra-Sosa et al., 2021). The fundamental step in the quantum processing pipeline is referred to as state preparation. It allows us to process data on quantum circuits before applying any specific quantum computation. Figure 28 shows an example of how quantum encoding is used to project data from the classical space to the quantum (Hilbert) space.
For tackling the data constraints in QML implementations, data encoding is a fundamental step for representing classical data as quantum states. Encoding layers largely influence the QML model expressivity since the data encoding strategy defines and drives the relevant QML parameters, e.g., the features the QML model can represent (Markidis, 2023). Knowing the various encoding techniques is necessary to choose the most suitable one for solving the particular QML task. Each encoding method has constraints due to the unstable quantum mechanical properties that hinder the complete data encapsulation in quantum representation. However, despite that, they successfully encode the information into quantum states and have large margins of improvement over time (LaRose & Coyle, 2020; Li et al., 2022; Weigold et al., 2021). Figure 30 shows a generic overview of the quantum state preparation stage. The process of embedding/encoding generates a quantum representation of the classical data governed by the selected encoding technique within the Hilbert space dimension.
@@ 5.4 Data Encoding
    The Basis Embedding encodes the binary feature vector into a basis state. It is primarily used when real numbers are mathematically manipulated in quantum algorithms. Such an encoding represents real numbers as binary numbers and transforms them into quantum states. However, since the binary features are not differentiable, the basis encoding does not allow gradient computations (Schuld & Petruccione, 2018).
@@
    In the Amplitude Embedding technique, classical data is encoded into the amplitudes of a quantum state. Besides defining the measurement probabilities, each quantum system’s wavefunction can be used to represent its amplitude as a data value. The main advantage of the amplitude embedding is that it can encode $n$ real values (or $n$ fixed-point precision approximation of real numbers) using $\mathcal{O}(\log{n})$ qubits, i.e., it has a logarithmic runtime dependency on the dataset size. Due to its efficiency, it is employed in many QML algorithms (Schuld & Petruccione, 2018).
@@
    The Angle Embedding is performed by applying rotations on the x-axis or y-axis using specialized quantum gates, called Pauli rotation gates, along with the values that must be encoded. If we apply angle embedding on a dataset, the number of rotations is the same as the number of features in the dataset (Schuld & Petruccione, 2018).
@@
    The Quantum Random Access Memory (QRAM) Encoding (Park et al., 2019) is a mechanism that allows accessing classically stored information in superposition by querying an index register. In other words, it enables direct access to the quantum data given the corresponding classical data value. By introducing a conditional rotation and branch selection procedure, we can feed the QRAM with the classical address of the required quantum data and get the qubit state at the output.
@@
    The Qsample Encoding associates a real amplitude vector with a classical discrete probability distribution. It can be considered a hybrid case of Basis and Amplitude Encoding because amplitudes represent the encoded information, but the features are encoded in the qubits (Schuld & Petruccione, 2018).
@@
    The Instantaneous Quantum Polynomial-time (IQP) Embedding (Havlíček et al., 2019) employs the so-called IQP circuit, which is a quantum circuit of a block of Hadamard gates, followed by a set of gates that are diagonal in the computational basis. Each diagonal gate is composed of a single-qubit RZ rotation gate, encoding the $n$ features, followed by a two-qubit ZZ entangler. The entangler, whose pattern can be customized, encodes the product between features.
    Table 2: Quantum datasets with key features.
@@
    QDataSet (Perrier et al., 2022) One-qubit and two-qubit systems evolving $14$ TB (compressed), $10,000$ samples each Aug.2021
    QM7, QM8, QM9 (Rupp et al., 2012) (Montavon et al., 2013) (Ramakrishnan et al., 2015) (Ramakrishnan et al., 2014) Organic molecules Between $7,165$ and $133,885$ molecules, with $7$-$9$ heavy atoms 2012-2015
    PennyLane Datasets (Bergholm et al., 2018) Organic molecules and spin $42$ geometries of $34$ molecules, $4$ spin models with $100$ configurations Apr.2022
    TensorFlow Quantum Data (Farhi & Neven, 2018) Image classification One-qubit per pixel Mar.2020
    Mendley Datasets for Quantum Circuit Mapping (Acampora et al., 2021) Circuit mapping problem as classification task $188,434$ random quantum circuits Sep.2021
    QFlow Lite (Zwolak et al., 2018) Charge state recognition $1,001$ $1,599$ measurements, $100\times 100$ $250\times 250$ pixel maps Sep.2018
    NTangled Datasets (Schatzki et al., 2021) Quantum states with multipartite entanglement $12,040$ training samples, $4,060$ test samples Nov.2021
@@
    The Quantum Approximate Optimization Algorithm (QAOA) Embedding (Farhi et al., 2014) encodes $n$ features into $m$ qubits, where $m>n$. It uses a layered trainable quantum circuit inspired by the QAOA Ansatz. The feature-encoding circuit associates features with the angles of RX rotations. The main advantage is that it supports gradient computations with respect to both the features and the weights.
@@
    The Hamiltonian Embedding scheme encodes features into the values of the Hamiltonian operator matrix (Schuld & Petruccione, 2018). The Hamiltonian operator of a system is a quantum mechanical operator that corresponds to the total energy of the system. Unlike previous methods that encode the data in the qubits’ states, the Hamiltonian embedding method encodes the data into the operator. Since it is highly correlated with physics, various physical QML algorithms, including VQE and QAOA, are developed using the Hamiltonian encoding method.
@@
    The Displacement Embedding encodes $n$ features into the displacement of amplitudes or phases of $m$ modes, where $n\leq m$. It is used in continuous-variable quantum computing models, where classical information is encoded in the displacement operator parameters (Markidis, 2023).
5.5 Quantum Datasets Collection